# Mining Legal Arguments to Study Judicial Formalism

### v1

@@ 2.4 Legal argument mining
Argument mining aims to automatically identify and extract arguments [lawrence_argument_2019]. Legal argument mining (LAM) focuses on the legal domain and legal arguments [habernal_mining_2024-1, zhang2022, MochalesMoens2011]. Recent research has applied legal argument mining to classify the traditional types of legal arguments [habernal_mining_2024-1, luders_classifying_2024-1]. This emerging field has been growing in recent years and increasingly intersects with empirical legal research. For recent advances, we refer to recent LAM papers [habernal_mining_2024-1, luders_classifying_2024-1] as well as a systematic literature review of the field [zhang2022].
@@
In sum, despite a century of debate about formalism and decades of strong claims about CEE judicial formalism, the field lacks both clear definitions and systematic evidence. Previous empirical studies remain small-scale and unreproducible, while legal argument mining has yet to be applied to CEE courts. We fill this gap by combining prominent works in legal theory with state-of-the-art NLP methods to test these longstanding claims at scale, as the next section explains.
3 Measuring formalism in judicial decisions
@@ 3.1 Five core tenets of CEE formalism
3.2 Measuring formalism through argument types
The key literature on formalism discusses legal argumentation; the three core tenets of CEE formalism concern arguments. We therefore begun with arguments.
@@ 3.2 Measuring formalism through argument types
On their own, the tenets are too abstract for empirical coding of concrete cases. For example, the first core tenet of CEE formalism (heavy reliance on text based arguments) requires a precise specification: what counts as a text-based argument, and how can we identify the linguistic argument is really present in an actual decision? Without such specification, empirically analyzing real world case law is challenging, if not impossible.
@@
To address this challenge, we developed a comprehensive annotation guidelines including an annotation scheme and instructions specifying what each type means and when to annotate it. The guidelines emerged iteratively: we began with leading argumentation taxonomies [alexy_argument_2010, maccormick_interpreting_2016, Walton_Macagno_Sartor_2021], incorporated German and Czech legal argumentation scholarship to represent literature from the Cetral European region [Mollers2020LegalMethods, Wintr2019Metody, LarenzCanaris_Methodenlehre_1995], applied and refined them on more than 200 Czech Supreme Court pilot decisions (SC and SAC), and then re-checked the adjustments against the same literature. We repeated these steps iteratively over several months before annotating the MADON dataset.
Our guidelines identified eight argument types, which we classify as formalistic or non-formalistic based on their relationship to the core tenets of CEE formalism:
Formalistic argument types
LIN
– Linguistic Interpretation
SI
– Systemic Interpretation (including Constitutional‑conforming and EU‑conforming interpretation)
CL
– Case Law
D
– Doctrine
Non-formalistic argument types
HI
– Historical Interpretation
PL
– Principles of Law and Values (including Constitutional Values, Rights, Principles and EU Values and Principles)
TI
– Teleological Interpretation
PC
– Practical Consequences
@@
For each argument type, our guideliness included a structured description that connects the theoretical concept to real‑world decisions. For most types, we (1) give a brief description, (2) list its subtypes, (3) note “trigger” phrases that usually signal its presence in a decision, (4) provide typical and borderline examples, (5) add comments and annotator tips and (6) include annotation chart consistent with [thalken_modeling_2023] to guide the annotation process. The creation of these guidelines itself represents an important contribution to the field. [^1] Appendix A contains the summary of the guideliness. Full version of the guideliness is accesible at https://github.com/trusthlt/madon/.
@@
We then applied the guideliness to annotate the below described MADON dataset of 272 decisions from Supreme Court and Supreme Administrative Court of Czechia. We used set of cases distinct from the pilot decisions to support reliability of the annotation.
The court decisions are structured into paragraphs. We reflected this structure and chose paragraph-level annotation over sentence-level coding for annotating arguments. Sentence-level coding produced very low Krippendorff unitized Alpha scores—annotators disagreed where exactly arguments start and end. From the perspective of our research question, we were not so much interested in whether a particular sentence belonged to an argument, but more about whether a particular argument is present in a certain paragraph. Thus, this decision should not negatively influence the results.
Each argument type would be counted only once per paragraph, even if multiple instances within one paragraph appeared. However, one paragraph could include two different argument types. For example, if a paragraph cited five previous judgments and referenced two explanatory notes, we would code it once for case law (CL – Case Law) and once for historical interpretation (HI – Historical Interpretation). We hypothesize that this approach reflects how legal theorists typically analyze argumentation: focusing on argument types rather than counting individual citations within one single sentence or paragraph.
We excluded rejected arguments from annotation. When courts discussed competing interpretations, we coded only the reasoning ultimately adopted. For example, if a court noted that historical interpretation supported outcome X while linguistic interpretation pointed to non-X, but the court then rejected linguistic interpretation and decided for X, we annotated only the historical interpretation.
@@ 3.3 Holistically labelling judgments as formalistic
Although arguments are crucial to determine prevailing reasoning style, formalism is not just about quantity of arguments. In our pilot studies, simply counting argument frequencies proved important, but insufficient. Some decisions dominated by formalistic arguments turned out non‑formalistic once assessed holistically—against the core tenets, the whole decision, its context, facts, and procedural history. Although quantification of arguments seems to be crucial prerequisite to study formalism and would likely suffice in most cases as our quantitative results show, it can also be misleading in some cases. There are three main reasons:
First, arguments differ in weight and form. A judgment may invoke a fundamental principle once or twice in a decisive paragraph, yet be packed with dozens of formalistic arguments in peripheral sections on procedure. Such a decision likely would be, in substance, non‑formalistic, but simple counting would suggest the opposite. Argument clarity also varies: some arguments are explicit; others are barely there and must be inferred from context. The problem deepens when courts cite statutory provisions that themselves embed non‑formalistic considerations (e.g., principles), blurring whether the reasoning is formalistic or not. And some “principles” are formalistic by design—think of the ”concentration principle,” which strictly bars new evidence in later stages of the proceedings. While arguments significantly matter, their quantification alone might therefore mislead.
Second, context matters. A Supreme Court decision critising lower courts for excessive formalism may still be filled with citations to statutory text and case law (formalist arguments) while ultimately pushing for a purposive, non-formalistic decision-making.
Third, the last two tenets, i.e., dismissal on formal grounds and insufficient reasoning, are not capturable by simply counting arguments.
To address this limitation, we annotated each decision with a binary label ‘formalistic/non-formalistic’. The expert decision is based on an holistic assessment of the entire case. The evaluations is complex and considers multiple contextual factors beyond argument counting: adherence to the five core tenets of formalism, the overall approach to legal interpretation, the weight and centrality of different arguments, whether the court critiques previous courts for formalistic reasoning, and the broader procedural and substantive context of the case.
@@ 3.4 Annotation process
The systematic annotation of Czech Supreme Courts’ decisions was conducted between July and September 2024, building on a pilot study from early 2024 that analyzed 160 decisions and helped refine the methodology. We employed four law students from the Faculty of Law, Charles University, as research assistants, that were lead by a PhD student in legal argumentation. All of them had a background in legal theory and argumentation.
The process had four phases: Introduction (Week 1), Training (Weeks 2–6, 60+ decisions), Coding (Weeks 7–12, 272 decisions with regular reviews), and Finalization (Week 13). The team met on weekly bases to discuss the annotations in the Introduction and Training phase. We used the INCEpTION [tubiblio106270] interface for accessible annotation. We followed methodological guidance for the legal domain and drawed mainly from content analysis literature [hall2008, Ovádek_Schroeder_Zglinski_2024] and recommended practices by braun_i_2023-1 for annotating legal NLP datasets. This approach involved several hundred hours of training and guideline development, with at least two independent annotators coding each decision in the MADON dataset. We measured inter-annotator agreement and report on the annotation process. In total, the annotation took 1,000+ hours, covered more than 450 decisions including pilot studies (272 after pilot), and produced the MADON dataset with 272 decisions consisting of 9,183 annotated paragraphs, including 1,913 legal arguments in total and 1,237 paragraphs with at least one argument.
@@ 3.4.1 Inter-annotator agreement and dataset curation
We measured inter-annotator agreement to asses reliability of our annotations. We calculated two different inter-annotator agreement metrics corresponding to our two types of labels: Cohen’s kappa for the holistic label and Krippendorff’s unitized Alpha for argument types.
Our holistic assessment achieved a Cohen’s kappa coefficient of 0.65 on the overall category formalistic/non-formalistic, indicating “substantial agreement” among annotators despite the complexity of the assessment [Landis_Koch_1977]. This success yields two significant findings for future empirical research. First, research assistants can be effectively trained to achieve sufficient inter-coder agreement on complex holistic assessments of formalistic/non-formalistic reasoning. Second, we found a strong correlation between decisions with higher frequencies of non-formalistic arguments and holistic nonformalistic assessments, suggesting that while the dual approach provides valuable validation, the relationship between argument types and overall formalism remains strong.
@@
For argument type annotation, we calculated Krippendorff’s unitized Alpha across two parts of the annotation process (in total 272 decisions), see Table 1. According to Krippendorff, sufficient agreement requires $\alpha\geq 0.800$, while data with agreement $0.800>\alpha\geq 0.67$ can be used to draw tentative conclusions [Krippendorff_2019]. Half of our categories meet the 0.67 threshold, while three categories (Linguistic Interpretation, Systematic Interpretation, and Practical Consequences) demonstrate significantly lower reliability. This pattern reflects the inherent complexity of legal document annotation. Annotation of legal documents is generally considered challenging, and inter-annotator agreement is often lower [braun_i_2023-1, vanDijck2024].
@@
Disagreements were resolved through “arbiter review” as suggested by [braun_i_2023-1] or a combination of arbiter review and “forced agreement” for complex cases – mainly the holistic label. The combination of arbiter review and forced agreement meant that annotators were required to discuss their disagreements and propose potential solutions, accompanied by justifications for their choices. The arbiter then reviewed both the suggested solutions and their justifications; when annotators agreed on how to solve initial disagreement but the arbiter disagreed with their consensus, the arbiter assigned the final label (i.e., the initial label of one of the annotators).
4 MADON dataset and first empirical findings
@@
The MADON [^2] MADON stands for Mining Arguments to Debunk Old Narratives dataset comprises 272 legal decisions from Supreme Court and Supreme Administrative Court, the only two Supreme Courts. [^3] Like Germany, Austria, Poland and many other countries in Europe, Czechia has a separate Constitutional Court “above” the Supreme Courts. It hears mainly constitutional complaints about alleged violations of fundamental rights. The Constitutional Courts is significantly distinct from the Supreme Court and the Supreme Administrative Court and is not included in our dataset.
@@ 4.1 Stratified sampling of documents for expert annotation
The MADON dataset spans the period from 1997 to 2024, with a focus on decisions made between 2003 and 2023. To ensure a representative sample, stratified sampling was employed, balancing cases across time periods, court agendas (civil, criminal, and administrative), and types of cases (procedural and on the merits). This method ensured proportional representation across the Supreme Court’s civil branch (122 cases), criminal branch (58 cases), and the Supreme Administrative Court (90 cases). Within these subpopulations, we drew random samples.
4.2 MADON statistics
@@
Figure 1: Documents in MADON are (a) long on average but rarely exceeding 6,000 tokens, and contain 7 arguments on average (b) with some outliers with almost 80 arguments.
@@
Figure 2: Without distinguishing between the two courts, the most prevalent argument type for the entire period 2003–2023 is Case Law (CL).
Argument types
The distribution of arguments reveals interesting patterns (see Figure 2). Linguistic interpretation, despite its theoretical importance for formalism, appears only 108 times in the whole MADON dataset. Another surprising aspect is the extremely high frequency of Case law (707 arguments) which we discuss below.
Table 2: Distribution of formalistic arguments through paragraphs and documents
Arguments
Linguistic
Interpretation
Systemic
Interpretation
Case Law
Doctrine
Paragraphs (frequency) 108 208 707 68
Documents (existence) 63 78 181 39
Table 3: Distribution of non-formalistic arguments through paragraphs and documents
Arguments
Historical
Interpretation
Principles of Law
and Values
Teleological
Interpretation
Practical
Consequences
Paragraphs (frequency) 58 311 315 138
Documents (existence) 22 117 124 68
Holistic label
The distribution of holistic labels differs across courts. The SC had 116 formalistic (64%) and 66 non-formalistic (36%) decisions for a total of 182 cases. The SAC had 45 formalistic (50%) and 45 non-formalistic (50%) decisions for a total of 90 cases. Combined, there were 161 formalistic (59%) and 111 non-formalistic (41%) decisions out of 272 total cases.
4.3 Key legal empirical findings from the MADON dataset
The MADON dataset allows us to get the first systematic picture of both Czech Supreme Courts’ reasoning practices. Contrary to the Tale of Two Courts discussed in Section 2.2, i.e., a Tale of formalistic SC and non-formalistic SAC, we found that during the SAC’s initial decade (2003-2013), both courts exhibited remarkably similar levels of formalism on most indicators. For instance, both courts had approximately two-thirds of their decisions classified as formalistic, with the SAC showing even slightly more formalism on some indicators.
However, the temporal trends in Figure 3 show a notable divergence which started around 2011 and continued throughout the second decade (2014-2023), where the SAC shifted towards non-formalistic reasoning. During this second period, the proportion of non-formalistic SAC decisions increased by 70 percent (Figure 3c), and the average number of non-formalistic arguments per decision rose by more than 100 percent (Figure 3b), while the SC’s practices remained relatively stable. This divergence is also evident in the ratio of non-formalistic to formalistic arguments (Figure 3a) and the decreasing share of SAC decisions with zero non-formalistic arguments (Figure 3d). This indicates that the Tale of Two Courts—the narrative about formalistic SC and non-formalistic SAC—became close to reality only in the later period, if ever.
(a) Non-formalistic/formalistic argument ratio
(b) Average non-formalistic arguments per decision
(c) Share of non-formalistic decisions
(d) Share of decisions with zero non-formalistic arguments
Figure 3: Temporal trends in judicial reasoning patterns: Supreme Court (SC) vs. Supreme Administrative Court (SAC), 2003-2023
Furthermore, the findings challenge stereotypes about CEE legal reasoning. As can be seen from Figure 2, both the SC and SAC use text-based arguments scarcely: linguistic interpretation (LIN) accounts for 5.7% of all arguments across both courts in 2003–2023. By contrast, non-formalistic arguments from legal principles and values (PL) and teleological interpretation (TI) are much more common, together making up 32.3% of all arguments. Case law (CL) is the most common type at 37.4%. Notably, case law was by far most dominant argument type already in 2006 at 29%; this indicates that while its use has increased over time (both in percentage share and absolute numbers), it was already the most frequent category of legal arguments at that earlier point in the study period.
These results demonstrate that Czech judicial reasoning, particularly its heavy reliance on case law and non-formalistic argumentation, diverges from traditional definitions of formalism prevalent in both CEE and U.S. contexts, where formalism is connected to textualism. We thus empirically refute the notion that Czech courts are text-bound and had been precedent-averse [manko_weeds_2013, matczak_eu_2015, Spamann_2024].
4.4 MADON dataset for NLP experiments

### v2

@@ 2.4 Legal argument mining
Argument mining aims to automatically identify and extract arguments [lawrence_argument_2019]. Legal argument mining (LAM) focuses on the legal domain and legal arguments [Habernal.et.al.2023.AILaw, zhang2022, MochalesMoens2011]. Recent research has applied legal argument mining to classify the traditional types of legal arguments [Habernal.et.al.2023.AILaw, luders_classifying_2024-1]. This emerging field has been growing in recent years and increasingly intersects with empirical legal research. For current advances, we refer to recent LAM papers [Habernal.et.al.2023.AILaw, luders_classifying_2024-1] as well as a systematic literature reviews of the field and related workshops [zhang2022, liepina2025legal, amelr2025preface].
@@
3 Operationalizing and annotating formalism
To train supervised classifiers for detecting formalism, we operationalize an abstract legal concept into concrete, annotatable labels through “core tenets of CEE formalism”.
@@ 3.1 Five core tenets of CEE formalism
Most of the tenets are not CEE specific. Although we extracted the core tenets from a thorough review of key CEE scholarship as we aim to empirically test existing claims about the CEE region, the main tenets appear shared beyond CEE. Similar criteria appear across very different regions and periods–from Weber’s sociology of law a century ago [weber_1922] and Schauer’s influential jurisprudential article in the 1980s [schauer_1988] to more recent computational work by stiglitz_h_historical_2024. In these works, reliance on the text of legal norms, focus on the most locally applicable rule rather than general values or principles, and avoidance of political, social, and economic reasoning often appear as important aspects of formalism.
3.2 Argument types
@@
The three core tenets of CEE formalism all concern legal arguments, but they are too abstract for direct annotation of real-world decisions. For example, the first tenet (heavy reliance on text-based arguments) requires specifying what counts as text-based argument and how to identify it in the judgments. Without such specification, empirical analysis remains impossible.
@@
To address this challenge, we developed annotation guidelines iteratively over several months. We started from leading argumentation taxonomies [alexy_argument_2010, maccormick_interpreting_2016, Walton_Macagno_Sartor_2021] and German and Czech legal scholarship [Mollers2020LegalMethods, Wintr2019Metody, LarenzCanaris_Methodenlehre_1995], which we refined on more than 200 pilot decisions before annotating the MADON dataset. The guidelines identify eight argument types, classified as formalistic or non-formalistic based on the core tenets of CEE formalism:
FORMALISTIC NON-FORMALISTIC
LIN Linguistic Interpretation HI Historical Interpretation
SI Systemic Interpretation PL Principles of Law/Values
CL Case Law TI Teleological Interpretation
D Doctrine PC Practical Consequences
The first three core tenets map directly to four of our argument types. Heavy reliance on text (Tenet 1) corresponds to Linguistic Interpretation; narrow focus on rules rather than principles (Tenet 2) corresponds to low frequency of Principles of Law and Values; and exclusion of external standards (Tenet 3) corresponds to low frequency of Teleological Interpretation and Practical Consequences.
We classify Systemic Interpretation, Case Law, and Doctrine as formalistic, and Historical Interpretation as non-formalistic, consistent with previous studies [matczak_constitutions_2010-2, matczak_eu_2015]. Our underlying reasoning is that Systemic Interpretation reflects the internal hierarchy of legal norms; Case Law relies on the authority of prior decisions and rules extracted from them; and Doctrine provides authoritative guidance, typically through commentaries that systematize case law.
Historical Interpretation is more complicated. In Germany, it could even be seen as formalistic due to its link to the “subjective theory” of interpretation, which is often contrasted with the more open “objective theory” focusing on the purpose of the law [Mollers2020LegalMethods]. In contrast, US textualists reject legislative history altogether [scalia1997matter] and US empirical studies accordingly connect legislative history with non-formalistic reasoning [stiglitz_h_historical_2024]. Similarly, we classify Historical Interpretation as non-formalistic. In this decision, we mainly reflect the CEE scholarship, which often considers legislative history as distinct from the more formalistic methods related to the text of the law [matczak_constitutions_2010-2, Wintr2019Metody]. In practical terms, this classification has minimal impact on our results, as Historical Interpretation is the least common argument type in our dataset amounting to only 3% of all arguments.
@@ D Doctrine PC Practical Consequences
For each argument type, our guidelines included a structured description that connects the theoretical concept to real‑world decisions. For most types, we (1) give a brief description, (2) list its subtypes, (3) note “trigger” phrases that usually signal its presence in a decision, (4) provide typical and borderline examples, (5) add comments and annotator tips and (6) include annotation chart consistent with thalken_modeling_2023 to guide the annotation process. The creation of these guidelines itself represents an important contribution to the field. [^2] Appendix A contains the summary of the guidelines. Full version of the guidelines is accessible at https://github.com/trusthlt/madon.
@@
We then applied the guidelines to annotate the below described MADON dataset of 272 decisions from Supreme Court and Supreme Administrative Court of Czechia. We used a set of cases distinct from the pilot decisions to support reliability of the annotation.
Court decisions are structured into paragraphs, and we followed this structure by annotating at the paragraph level rather than sentence level. Our interest was in whether a given argument type is present in a paragraph, not in its exact boundaries. Each argument type was counted once per paragraph; each paragraph could thus include between zero and eight arguments. We annotated only the reasoning ultimately adopted and we excluded arguments the court explicitly rejected similarly to Krishnakumar2020.
@@ 3.3 Holistically labelling judgments as formalistic
Although arguments are important, we found we should not reduce formalism to simple quantification of arguments. In our pilot studies, some decisions dominated by formalistic arguments turned out non-formalistic once assessed holistically. Three reasons explain why counts alone can mislead. First, arguments differ in weight (some arguments are more important for the outcome while others are peripheral); context matters (a decision criticizing lower courts for excessive formalism may still cite mostly statutory text); and the last two tenets, i.e., dismissal on formal grounds and insufficient justification, are not capturable by simply counting arguments.
We therefore annotated each decision with a binary holistic label (formalistic/non-formalistic). The expert decision is based on a complex assessment of the entire case. Concretely, the annotators assessed each decision against all five core tenets. The annotators also considered the frequency, weight and centrality of individual arguments—for instance, whether a single decisive reference to a legal principle determined the result and outweighed numerous peripheral citations to statutory text, as well as other contextual and factual matters, like whether the court itself critiqued lower courts for excessive formalism. The label thus reflects an expert judgment about the decision as a whole. Our annotation guidelines include the instructions and annotation chart used for annotation. [^3] https://github.com/trusthlt/madon
@@ 3.4 Annotation process
The systematic annotation of Czech Supreme Courts’ decisions was conducted between July and September 2024, building on a pilot study from early 2024 that analyzed 160 decisions and helped refine the methodology. We employed four law students from the Faculty of Law, Charles University, as research assistants, who were led by a PhD student in legal argumentation. All of them had a background in legal theory and argumentation.
The process had four phases: Introduction (Week 1), Training (Weeks 2–6, 60+ decisions), Coding (Weeks 7–12, 272 decisions with regular reviews), and Finalization (Week 13). The team met on a weekly basis to discuss the annotations in the Introduction and Training phase. We used the INCEpTION [tubiblio106270] interface for accessible annotation. We followed methodological guidance for the legal domain and drew mainly from content analysis literature [hall2008, Ovádek_Schroeder_Zglinski_2024] and recommended practices by braun_i_2023-1 for annotating legal NLP datasets. This approach involved several hundred hours of training and guideline development, with at least two independent annotators coding each decision in the MADON dataset. We measured inter-annotator agreement and report on the annotation process. In total, the annotation took 1,000+ hours, covered more than 450 decisions including pilot studies (272 after pilot), and produced the MADON dataset with 272 decisions consisting of 9,183 annotated paragraphs, including 1,913 legal arguments in total and 1,237 paragraphs with at least one argument.
@@ 3.4.1 Inter-annotator agreement and dataset curation
To measure reliability, we calculated two different inter-annotator agreement metrics corresponding to our two types of labels: Cohen’s kappa for the holistic label and Krippendorff’s unitized alpha for argument types.
Our holistic assessment achieved a Cohen’s kappa coefficient of 0.65 on the overall category formalistic/non-formalistic, indicating “substantial agreement” among annotators despite the complexity of the assessment [Landis_Koch_1977]. This promising result yields two significant findings for future empirical research. First, research assistants can be effectively trained to achieve sufficient inter-coder agreement on complex holistic assessments of formalistic/non-formalistic reasoning. Second, we found a strong correlation between decisions with higher frequencies of non-formalistic arguments and holistic non-formalistic assessments; we use this finding to build original NLP architecture that predicts the holistic label based on the arguments.
@@
For argument type annotation, we calculated Krippendorff’s unitized alpha across two parts of the annotation process (in total 272 decisions), see Table 1. According to Krippendorff, sufficient agreement requires $\alpha\geq 0.800$, while data with agreement $0.800>\alpha\geq 0.667$ can be used to draw tentative conclusions [Krippendorff_2019]. Half of our categories meet the 0.667 threshold, while three categories (Linguistic Interpretation, Systemic Interpretation, and Practical Consequences) demonstrate significantly lower reliability.
This pattern reflects the inherent complexity of legal document annotation. Annotation of legal documents is generally considered challenging, and inter-annotator agreement is often lower [braun_i_2023-1, vanDijck2024]. Existing literature identifies three main sources of annotator disagreement: annotators themselves, annotation guidelines, and the concepts of interest [Ovádek_Schroeder_Zglinski_2024].
We believe the problem with the three categories (Linguistic Interpretation, Systemic Interpretation, and Practical Consequences) lies primarily with the concepts. Notably, most disagreements were reasonable, i.e., good reasons existed for either annotation, and clear mistakes or deviations from guidelines were rare.
For the three low-agreement argument types, we were unable to resolve the ambiguity despite extensive efforts in our pilot studies. We tried repeated revision rounds [vanDijck2024], better definitions, label refinement (e.g., merging constitutional conforming interpretation into systemic interpretation), additional examples, and flowcharts [thalken_modeling_2023]. These techniques helped but produced only modest improvements for the three argument types. The disagreements appear to reflect a deeper problem, namely that legal theory remains abstract and imprecise when applied to actual decisions. Most real-world cases are not, by definition, the prototypical examples from the textbooks, and the argument types themselves prove contentious and unstable in practice, even though they are widely used and broadly understood at the theoretical level. Adding more examples or decision rules did not resolve this, because the problem seems to lie with the concepts, not the guidelines.
We retain the three low-agreement categories despite the lower agreement. We do so for two main reasons. First, excluding labels due to low inter-annotator agreement is not standard practice in legal NLP; many studies either do not employ multiple annotators or do not report agreement at all [braun_i_2023-1]. Second reason is appropriate curation. We followed the recommended curation strategy; every disagreement was resolved by an arbiter, a co-author with legal background specializing in legal argumentation. The arbiter evaluated each case against the annotation guidelines and existing theories of legal argumentation before assigning one of the labels proposed by the original annotators. Every label in the dataset was thus approved by at least two people, with contested labels decided by a more senior domain expert. The annotator disagreement thus served a twofold goal; it first helped refine the guidelines during pilot studies, and later flagged for senior review those ambiguous cases that include something potentially relevant for our project. While lower agreement introduces some necessary noise, this curation strategy reduces its negative impact on the validity and consistency of the dataset. We follow the best practices and publish not only the curated data, but also all labels from all annotators.
@@
For the holistic label, we employed a combination of arbiter review and “forced agreement” [braun_i_2023-1]. This meant that annotators were required to discuss their disagreements and propose potential solutions, accompanied by justifications for their choices. The arbiter then reviewed both the suggested solutions and their justifications; when annotators agreed on how to solve initial disagreement but the arbiter disagreed with their consensus, the arbiter assigned the final label (i.e., the initial label of one of the annotators).
4 MADON dataset
@@
The MADON [^4] MADON stands for Mining Arguments to Debunk Old Narratives is the annotated dataset resulting from the process described in Section 3 and comprises 272 legal decisions from the Supreme Court and the Supreme Administrative Court, the only two Supreme Courts. [^5] Like Germany, Austria, Poland and many other countries in Europe, Czechia has a separate Constitutional Court “above” the Supreme Courts. It hears mainly constitutional complaints about alleged violations of fundamental rights. The Constitutional Court is significantly distinct from the Supreme Court and the Supreme Administrative Court and is not included in our dataset.
@@ 4.1 Stratified sampling of documents for expert annotation
The MADON dataset spans the period from 1997 to 2024, with a focus on decisions made between 2003 and 2023. We employed stratified sampling to ensure a representative sample. The stratified sample reflects time periods, court agendas (civil, criminal, and administrative), and types of cases (procedural and on the merits). This method ensured proportional representation across the Supreme Court’s civil branch (122 cases), criminal branch (58 cases), and the Supreme Administrative Court (90 cases). Within these strata, we drew random samples.
4.2 MADON statistics and first empirical findings
@@
Figure 1: Documents in the MADON dataset are (a) long on average but rarely exceeding 6,000 tokens, and contain 7 arguments on average (b) with some outliers with almost 80 arguments.
@@
Figure 2: Without distinguishing between the two courts, the most prevalent argument type for the entire period 1997–2024 is Case Law (CL).
The distribution of arguments reveals surprising patterns (see Figure 2). Linguistic interpretation (LIN), despite its theoretical centrality to formalism, accounts for only 5.7% of all arguments across both courts in 1997–2024 (108 occurrences). By contrast, two non-formalistic arguments, namely legal principles and values (PL) and teleological interpretation (TI), together make up 32.3%. Most striking is the dominance of case law (CL), which at 37.4% (707 occurrences) is the single most common argument type. These patterns challenge prevailing stereotypes: the SC is not text-bound, nor does it differ markedly from the SAC in argumentation profile. More broadly, Czech courts’ scarce use of explicit text-based reasoning and frequent reliance on purposive and value-based arguments diverges from formalism as understood in both CEE and U.S. contexts, where it is closely tied to textualism. We thus empirically challenge the claim that Czech courts are text-bound and had been precedent-averse [manko_weeds_2013, matczak_eu_2015, Spamann_2024].
While the argumentation profiles are similar at SC and SAC, the distribution of holistic formalism/non-formalism labels differs across courts. The SC had 116 formalistic (64%) and 66 non-formalistic (36%) decisions for a total of 182 cases. The SAC had 45 formalistic (50%) and 45 non-formalistic (50%) decisions for a total of 90 cases. Non-formalistic decisions are more frequent especially in the later period (2014-2024). This indicates that the alleged differences between the SC and SAC, i.e., the claim about formalistic SC and non-formalistic SAC, became closer to reality in the later period, where SAC moved toward more non-formalistic decisionmaking. These findings are described in more detail in our complementary work [koref2025]. The findings are based on the MADON dataset and await verification through our downstream large-scale analysis. In this paper, we use the MADON dataset primarily for the NLP experiments described in the following section.
Table 2: Distribution of argument types across paragraphs and documents
Type Argument Paragraphs Documents
Formalistic Linguistic Interpretation (LIN) 108 63
Systemic Interpretation (SI) 208 78
Case Law (CL) 707 181
Doctrine (D) 68 39
Non-formalistic Historical Interpretation (HI) 58 22
Principles of Law and Values (PL) 311 117
Teleological Interpretation (TI) 315 124
Practical Consequences (PC) 138 68
5 Experiments
5.1 Experiment setup and tasks
We decompose the problem into three tasks of increasing complexity, following standard practice in argument mining [eger-etal-2017-neural]:
  • Task 1 (argument presence detection) classifies paragraphs as argumentative or non-argumentative.
  • Task 2 (argument type classification) assigns up to eight argument type labels per paragraph in a multilabel setting.
  • Task 3 (holistic formalism classification) classifies entire decisions as formalistic or non-formalistic.
5.2 Data for NLP experiments